OpenStudy (anonymous):

Use the intercepts to graph the equation: x+2y=4 How would I go about doing this

OpenStudy (anonymous):

y= (-x+4)/2

OpenStudy (anonymous):

\[y=-\frac{1}{2}x+2\]

OpenStudy (anonymous):

alsamixer, when you are dont here, can you look at something for me?

OpenStudy (anonymous):

nevermind lol

OpenStudy (anonymous):

so would my graph look lie this when im done?

OpenStudy (anonymous):

no, y intercept = 2 x intercept = 4

OpenStudy (anonymous):

It's clear that \[\frac{x}{\text{x-intercept}}+\frac{y}{\text{y-intercept}}=1\] is the equation of a straight line. Because, set x = 0, you get the y = y-intercept and vice versa. Now, your problem is simple, just divide by the right hand side to get, \[\frac{x}{4}+\frac{y}{2}=1\] intercepts are 4 and 2 :)

OpenStudy (anonymous):

exactly, the graph is wrong!

OpenStudy (anonymous):

oh yeah.. lol guess it shows that i haven't done graphing for about 8 years :P

OpenStudy (anonymous):

it shud be like this the red line is the one u need

OpenStudy (anonymous):

sunu000 what software you use draw the graph?

OpenStudy (anonymous):

Kmplot...i am on my ubuntu linux now..:p.. i have a dual boot ... u can use microsoft math for windows i think :)

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

Microsft draw graph?

OpenStudy (anonymous):

microsoft math draws graphs i think...u just put in a function and it draws the graph...

OpenStudy (anonymous):

@arman: I solved one of your posted questions as well :)

OpenStudy (anonymous):

put x=0 you get y=2 and putting y=0 you get x=4 so along the axis mark the point A(4,0) and along y-axis mark the point B(0,2) and join AB...then AB will be the reqd line